OpenStudy (anonymous):

Find an equation for the nth term of a geometric sequence where the second and fifth terms are -2 and 16, respectively.

OpenStudy (anonymous):

@Loser66

OpenStudy (anonymous):

@amistre64 @hartnn

OpenStudy (cwrw238):

nth term = a1* r^(n-1) so 5th term = a1*r^4 and 2nd term = a1*r this should help r = common ratio which is the value you multiply each term to get the next term

OpenStudy (anonymous):

Is it an = 1 * (-2)^n-1?

OpenStudy (cwrw238):

r^3 = 16 / -2 = -8 so r = -2

OpenStudy (cwrw238):

so a1 = -2/-2 = 1

OpenStudy (anonymous):

a1 must = 1, r = -2 then 1*(-2)^n-1, thanks man!

OpenStudy (cwrw238):

you r right

OpenStudy (cwrw238):

yw - actually i'm a woman

OpenStudy (anonymous):

Oh lol ok, won't happen again :)

OpenStudy (cwrw238):

maybe a write like a man! - that's ok

OpenStudy (anonymous):

Nononono! I wasnt saying that lol!! The internet is tricky sometimes! :P